# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 5. Spectral average of Rankin-Selberg integrals
The space $S_{\varrho}({\bf K}_{0}(N))$ is endowed with the Hermitian inner product associated to the norm $\int_{{\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})}(\varphi(g)\mid\varphi(g))_{\varrho}\,{\rm{d}}g$ $(\varphi\in S_{\varrho}({\bf K}_{0}(N))$. Let ${\mathscr{H}}(\mathsf{G}({\mathbb{Q}}_{p})\mathbin{/\mkern-6.0mu/}{\bf K}_{p})$ be the Hecke algebra for $(\mathsf{G}({\mathbb{Q}}_{p}),{\bf K}_{p})$ for $p<\infty$. For any finite set $S$ of primes $p$ that is prime to $N$, define ${\mathscr{H}}_{S}:=\otimes_{p\in S}{\mathscr{H}}(\mathsf{G}({\mathbb{Q}}_{p})\mathbin{/\mkern-6.0mu/}{\bf K}_{p})$. The ${\mathbb{C}}$-algebra ${\mathscr{H}}_{S}$ acts on the finite dimensiona Hilbert space $S_{\varrho}({\bf K}_{0}(N))$ normally by
$[R(f_{S})\varphi](g)=\int_{\mathsf{G}({\mathbb{Q}}_{S})}\varphi(gx_{S})f_{S}(x_{S})\,{\rm{d}}x_{S},\quad g\in\mathsf{G}({\mathbb{A}}),\quad f_{S}\in{\mathscr{H}}_{S},\,\varphi\in S_{l}({\bf K}_{0}(N)).$
where $\mathsf{G}({\mathbb{Q}}_{S}):=\prod_{p\in S}\mathsf{G}({\mathbb{Q}}_{p})$. Let $\mu$ and $M$ be as in §4.3. We define the Schwarts-Bruhat function $\phi\in\mathcal{S}({\mathbb{A}}_{E}^{2})$ associated $\mu$ as
$\phi=\prod_{p{\leqslant}\infty}\phi_{p},\quad\phi_{p}(x,y)=\begin{cases}\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{E_{p}}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{E_{p}}}(y)\quad&(p<\infty,p\nmid M)\\ \operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{p^{e}\mathfrak{o}_{E_{p}}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{1+p^{e}\mathfrak{o}_{E_{p}}}(y)&(p<\infty,p^{e}\parallel M,e{\geqslant}1),\end{cases}$
@@
(5.2) $b_{p}=\begin{cases}1_{4}\quad&(p<\infty,p\nmid NM),\\ \eta_{p}&(p<\infty,p\mid N),\\ b_{p}^{M}:=\left[\begin{smallmatrix}p^{e}1_{2}&T_{\theta}^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]&(p<\infty,p^{e}\parallel M,e{\geqslant}1),\\ b_{\mathbb{R}}^{\theta}&(p=\infty)\end{cases}$
@@
where $\widehat{f_{S}}(\pi_{S})$ is the spherical Fourier transforms of $f_{S}$ on $\pi_{S}:=\otimes_{p\in S}\pi_{p}$, which is defined as the eigenvalue of the operator $\pi_{S}(f_{S})$ on the ${\bf K}_{S}=\prod_{p\in S}{\bf K}_{p}$-fixed vectors $\pi_{S}^{{\bf K}_{S}}\cong{\mathbb{C}}$. For $\bullet\in\{{\rm T},{\rm G},{\rm Y},{\rm SK}\}$ and $\ast\in\{{\rm new},{\rm old}\}$, we define ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet}$, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\ast}$, and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet,\ast}$ to be ${\mathbb{I}}^{(s)}(X,f_{S})$ with $X=\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet}$, $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\ast}$ and $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet,\ast}$, respectively. Then, due to (4.18), the average (5.1) is rearranged as
@@ 5.1. A construction of orthonormal basis
Let $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}$ be an element of the set $\Pi_{\rm csup}^{(T_{\theta},\Lambda)}(\lambda,N)$ (see §4.3). We fix a $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ of $\pi$ (see §4.4) once and for all. Set $\varphi_{\pi}^{0}(g):=(\varphi_{\pi,\varrho}^{0}(g)\mid\mathsf{v}_{\varrho^{0}})_{\varrho}$, which is an element of $V_{\pi}^{{\bf K}_{0}(N)}$. Since $\pi$ is a unitary representation, all of its factors $\pi_{p}$ are unitarizable. For each $p<\infty$, we can uniquely fix a $\mathsf{G}({\mathbb{Q}}_{p})$-invariant inner product $(\cdot|\cdot)_{p}$ on $V_{\pi_{p}}$ by demanding $(\xi_{p}|\xi_{p})_{p}=1$; for $p=\infty$, we fix a $\mathsf{G}({\mathbb{R}})$-invarinat inner product so that its pull back to $V_{\varrho}\hookrightarrow\pi_{\infty}$ (cf. §4.2(i)) coincides with the inner product $(\cdot|\cdot)_{\varrho}$ of $V_{\varrho}$. Let $\varphi_{\pi}^{0}$ be the global new form attached to $\{\xi_{p}\}_{p<\infty}$ (see §4.4). Then,
(5.5) $\frac{(\varphi|\varphi)_{L^{2}}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}=(v|v)_{\varrho}\prod_{p<\infty}(v_{p}|v_{p})_{p}$
@@
(5.6) $(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}^{-1/2}\times\otimes_{p<\infty}v_{p},\quad v_{p}\in{\mathcal{B}}(\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p})),$
@@ 5.2. Computation of the average
(5.7) $\langle E(\phi,s,\Lambda,\mu),R(b)\overline{\varphi_{\mathsf{v}_{\varrho}^{0}}}\rangle$ $=(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}^{-1/2}\,w_{D}^{-1}\,\overline{R(\varphi^{0}_{\pi},E,\Lambda)}\,\widehat{L}\left(s+\tfrac{1}{2},\pi,\mu\right)$
@@
Let $(\lambda=(l_{1},l_{2}),N)$, $\mu$, $\tilde{\mu}$ and $M$ be as in §4.3. Let $S$ be a finite set of prime numbers relatively prime to $DMN$. Then, for any $f_{S}=\otimes_{p\in S}f_{p}\in{\mathscr{H}}_{S}$,
(5.11) ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})$ $=\frac{2^{-2}|D|^{\frac{1}{2}(3-\frac{l_{1}+l_{2}}{2})}e^{-2\pi\sqrt{|D|}}}{w_{D}^{2}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\mid N}(1+p^{-2})^{-1}$
@@
with ${\mathfrak{t}}(\pi,\mu)=\prod_{p\mid N}{\mathfrak{t}}(\pi_{p},\mu_{p})$ with
@@
To refine (5.11) by considering ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm T},{\rm new}}$ and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{{\rm SK},{\rm new}}$, we refer to the explicit formula of the quantity $\frac{|R(\varphi^{0}_{\pi},E,\Lambda)|^{2}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}$ first proved by Dickson, Pitale, Saha, and Schmidt ([7, Theorems 1.13 and 3.13]) when $l_{1}=l_{2}$ under the refined GGP conjecture for Bessel periods posed by Liu([23]). Owing to a theorem by Furusawa and Morimoto [11, Theorem 8.1], the formula in [7, Theorems 1.13 and 3.13] is extended to vector-valued forms unconditionally so that it can applied to (5.11) to yield the following theorem. Note that the polynomial $Q_{S,\varrho}$ (with $S=T_{\theta}$) in [10, (8.2.16)] equals $(-1)^{\frac{l_{1}-l_{2}}{2}}(\frac{2}{\sqrt{|D|}})^{\frac{l_{1}+l_{2}}{2}}\mathsf{v}_{\varrho}^{\theta}$, our $R(\varphi,E,\Lambda)$ equals the quantity $w_{D}(\frac{\sqrt{|D|}}{2})^{-\frac{l_{1}+l_{2}}{2}}\,\mathcal{B}_{\Lambda}(\varphi;E)/(Q_{S,\varrho},Q_{S,\varrho})_{l_{1}-l_{2}}$ defined by [10, (8.2.17)], and $(Q_{S,\varrho},Q_{S,\varrho})_{l_{1}-l_{2}}=(\frac{|D|}{4})^{-\frac{l_{1}+l_{2}}{2}}(\mathsf{v}_{\varrho}^{0},\mathsf{v}_{\varrho}^{0})_{l_{1}-l_{2}}=(\frac{|D|}{4})^{-\frac{l_{1}+l_{2}}{2}}(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}$.
@@
Let the notations and the assumptions be as in Theorem 5.2. We have
@@
    ×$\times M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})$
@@
    where $S(N)$ is the set of all the prime divisors of $N$, and ${\mathcal{A}I}(\Lambda^{-1})$ is the automorphic representation of ${\bf GL}_{2}({\mathbb{A}})$ defined by the automorphic induction of $\Lambda^{-1}$ from ${\mathbb{A}}_{E}^{\times}$.
@@
    ×$\times M^{s-6}\zeta_{M}(1)\zeta_{M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})$
@@
The equation (i) and (ii) is a direct corollary to Theorem 5.3 and [10, Theorem 8.1](for the scalar case we refer to [7, Theorems 1.13 and 3.14]; we also note the formula
@@
where $\pi=Y(\pi_{1},\pi_{2})$. This obtained immediately by comparing the local factors. It is noted in [7, p.296] that each local representation $\pi_{p}$ of $\pi\cong\otimes_{p}\pi_{p}\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm Y,new}$ for $p\mid N$ is of type VIb. Finally, we check the statement (iii). As quoted in [7, Theorem 3.11], the $(T_{\theta},\Lambda)$-Bessel periods of the Saito-Kurokawa lifts are zero unless $\Lambda$ is trivial due to Qiu. By comparing the local $L$-functions ([39, Theorem 5.2(ii)], cf. [29, Theorem 3.1]),
@@
By (5.4), we have that ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})-{\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm G,new}$ is the sum of ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm Y,new}$, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm SK,new}$ and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm old}$. Our main focus is their values at $s=0$. For those, we have the following upper bounds in the level aspect.
@@
    Let $\lambda$ be fixed as before and $N$ an inert prime in $E/{\mathbb{Q}}$. Then, there exist constants $C_{D,\lambda,\mu,f_{S}},C_{D,\lambda,\mu,f_{S}}^{\prime}>0$ depending only on $D$, $l$, $\mu$, and $f_{S}$ such that
    (5.14) $|{\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm old}}|<C_{D,\lambda,\mu,f_{S}}N^{-8}.$
    (5.15) $|{\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm Y,new}}|<C_{D,\lambda,\mu,f_{S}}^{\prime}N^{-\frac{3}{2}}.$
@@
    Let $N$ be $1$ or an inert prime in $E/{\mathbb{Q}}$. When $\mu\neq{\bf 1}$, ${\mathbb{I}}^{(0)}(l,N,f_{S})^{{\rm SK,new}}=0.$ When $\mu={\bf 1}$, there exists a constant $C_{f_{S}}>0$ depending only on $f_{S}$ such that
    (5.16) $|{\mathbb{I}}^{(0)}((l,l),N,f_{S})^{{\rm SK,new}}|<C_{f_{S}}|D|^{\frac{l}{2}}l^{\frac{1}{2}}N^{-\frac{3}{2}}.$
@@
Because $N$ is a prime, one can check that the summation range of ${\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm old}}$ is $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{{\rm old}}=\Pi_{\rm cusp}^{(T_{\theta},\chi)}(\lambda,1)$, which is independent of $N$. From the formula of ${\mathfrak{t}}(\pi_{p},\mu_{p})$ in (5.12) for $p=N$, combined with the temperedness of $\pi_{N}$ due to [48] and the matrix of $T_{1,0}$ in table (6.49), we get ${\mathfrak{t}}(\pi_{N},\mu_{N})=O(N^{-4})$. This and the equation $[{\bf K}_{\bf f}:{\bf K}_{0}(N)]=N^{3}(1+N^{-2})(1+N^{-1})$ for prime $N$ yield the bound (5.14).

### v2

@@ 5. Spectral average of Rankin-Selberg integrals
The space $S_{\varrho}({\bf K}_{0}(N))$ is endowed with the Hermitian inner product associated to the norm $\int_{{\mathsf{Z}}({\mathbb{A}})\mathsf{G}({\mathbb{Q}})\backslash\mathsf{G}({\mathbb{A}})}(\varphi(g)\mid\varphi(g))_{\varrho}\,{\rm{d}}g$ $(\varphi\in S_{\varrho}({\bf K}_{0}(N))$. Let ${\mathscr{H}}(\mathsf{G}({\mathbb{Q}}_{p})\mathbin{/\mkern-6.0mu/}{\bf K}_{p})$ be the Hecke algebra for $(\mathsf{G}({\mathbb{Q}}_{p}),{\bf K}_{p})$ for $p<\infty$. For any finite set $S$ of primes $p$ that is prime to $N$, define ${\mathscr{H}}_{S}:=\otimes_{p\in S}{\mathscr{H}}(\mathsf{G}({\mathbb{Q}}_{p})\mathbin{/\mkern-6.0mu/}{\bf K}_{p})$. The ${\mathbb{C}}$-algebra ${\mathscr{H}}_{S}$ acts on the finite dimensional Hilbert space $S_{\varrho}({\bf K}_{0}(N))$ normally by
$[R(f_{S})\varphi](g)=\int_{\mathsf{G}({\mathbb{Q}}_{S})}\varphi(gx_{S})f_{S}(x_{S})\,{\rm{d}}x_{S},\quad g\in\mathsf{G}({\mathbb{A}}),\quad f_{S}\in{\mathscr{H}}_{S},\,\varphi\in S_{l}({\bf K}_{0}(N)),$
where $\mathsf{G}({\mathbb{Q}}_{S}):=\prod_{p\in S}\mathsf{G}({\mathbb{Q}}_{p})$. Let $\mu$ and $M$ be as in §4.3. We define the Schwartz-Bruhat function $\phi\in\mathcal{S}({\mathbb{A}}_{E}^{2})$ associated with $\mu$ as
$\phi=\prod_{p{\leqslant}\infty}\phi_{p},\quad\phi_{p}(x,y)=\begin{cases}\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{E_{p}}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{\mathfrak{o}_{E_{p}}}(y)\quad&(p<\infty,p\nmid M{\mathcal{M}})\\ \operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{p^{e}\mathfrak{o}_{E_{p}}}(x)\operatorname{\mbox{{\rm 1}}\mbox{{\rm l}}}_{1+p^{e}\mathfrak{o}_{E_{p}}}(y)&(p<\infty,p^{e}\parallel M{\mathcal{M}},e{\geqslant}1),\end{cases}$
@@
(5.2) $b_{p}=\begin{cases}1_{4}\quad&(p<\infty,p\nmid NM),\\ \eta_{p}&(p<\infty,p\mid N),\\ b_{p}^{M}:=\left[\begin{smallmatrix}p^{e}1_{2}&T_{\theta}^{\dagger}\\ 0&1_{2}\end{smallmatrix}\right]&(p<\infty,p^{e}\parallel M,e{\geqslant}1),\\ b_{p}^{{\mathcal{M}}}:=\left[\begin{smallmatrix}p^{2}&&&\\ &p&&\\ &&1&\\ &&&p\end{smallmatrix}\right]&(p\mid{\mathcal{M}}),\\ b_{\mathbb{R}}^{\theta}&(p=\infty)\end{cases}$
@@
where $\widehat{f_{S}}(\pi_{S})$ is the spherical Fourier transforms of $f_{S}$ at $\pi_{S}:=\otimes_{p\in S}\pi_{p}$, which is defined as the eigenvalue of the operator $\pi_{S}(f_{S})$ on the ${\bf K}_{S}=\prod_{p\in S}{\bf K}_{p}$-fixed vectors $\pi_{S}^{{\bf K}_{S}}\cong{\mathbb{C}}$. For $\bullet\in\{{\rm T},{\rm G},{\rm Y},{\rm SK}\}$ and $\ast\in\{{\rm new},{\rm old}\}$, we define ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet}$, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\ast}$, and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet,\ast}$ to be ${\mathbb{I}}^{(s)}(X,f_{S})$ with $X=\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet}$, $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\ast}$ and $\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\bullet,\ast}$, respectively. Then, due to (4.18), the average (5.1) has the expression:
@@ 5.1. A construction of orthonormal basis
Let $\pi\cong\otimes_{p\leqslant\infty}\pi_{p}$ be an element of the set $\Pi_{\rm csup}^{(T_{\theta},\Lambda)}(\lambda,N)$ (see §4.3). We fix a $(T_{\theta},\Lambda)$-Bessel data $\{(\ell_{p},\xi_{p})\}_{p<\infty}$ of $\pi$ (see §4.4) once and for all. Set $\varphi_{\pi}^{0}(g):=(\varphi_{\pi,\varrho}^{0}(g)\mid\mathsf{v}_{\varrho}^{0})_{\varrho}$, which is an element of $V_{\pi}^{{\bf K}_{0}(N)}$. Since $\pi$ is a unitary representation, all of its factors $\pi_{p}$ are unitarizable. For each $p<\infty$, we can uniquely fix a $\mathsf{G}({\mathbb{Q}}_{p})$-invariant inner product $(\cdot\mid\cdot)_{p}$ on $V_{\pi_{p}}$ by demanding $(\xi_{p}\mid\xi_{p})_{p}=1$; for $p=\infty$, we fix a $\mathsf{G}({\mathbb{R}})$-invarinat inner product so that its pull back to $V_{\varrho}\hookrightarrow\pi_{\infty}$ (cf. §4.2(i)) coincides with the inner product $(\cdot\mid\cdot)_{\varrho}$ of $V_{\varrho}$. Let $\varphi_{\pi}^{0}$ be the global new form attached to $\{\xi_{p}\}_{p<\infty}$ (see §4.4). Then,
(5.5) $\frac{(\varphi\mid\varphi)_{L^{2}}}{(\varphi_{\pi}^{0}\mid\varphi_{\pi}^{0})_{L^{2}}}=(v|v)_{\varrho}\prod_{p<\infty}(v_{p}|v_{p})_{p}$
@@
(5.6) $(\varphi_{\pi}^{0}\mid\varphi_{\pi}^{0})_{L^{2}}^{-1/2}\times\otimes_{p<\infty}v_{p},\quad v_{p}\in{\mathcal{B}}(\pi_{p},{\bf K}_{0}(N{\mathbb{Z}}_{p})),$
@@ 5.2. Computation of the average
(5.7) $\langle E(\phi,s,\Lambda,\mu),R(b)\overline{\varphi_{\mathsf{v}_{\varrho}^{0}}}\rangle$ $=(\varphi_{\pi}^{0}\mid\varphi_{\pi}^{0})_{L^{2}}^{-1/2}\,w_{D}^{-1}\,\overline{R(\varphi^{0}_{\pi},E,\Lambda)}\,\widehat{L}\left(s+\tfrac{1}{2},\pi,\mu\right)$
@@
Let $(\lambda=(l_{1},l_{2}),N)$, $\mu$, $\tilde{\mu}$ and $M$ be as in §4.3. Let $S$ be a finite set of prime numbers relatively prime to $DMN$. Then, for any $f_{S}=\otimes_{p\in S}f_{p}\in{\mathscr{H}}_{S}$, the value ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})$ equals
(5.11) $\frac{2^{-2}|D|^{\frac{1}{2}(3-\frac{l_{1}+l_{2}}{2})}e^{-2\pi\sqrt{|D|}}}{w_{D}^{2}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]}{\mathcal{M}}^{-4}M^{s-6}\zeta_{M}(1)\zeta_{M{\mathcal{M}}}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\mid N}(1+p^{-2})^{-1}$
@@
with ${\mathfrak{t}}(\pi,\mu)=\prod_{p\mid N}{\mathfrak{t}}(\pi_{p},\mu_{p})$,
@@
To refine (5.11) by considering ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\bullet,{\rm new}}$ ($\bullet\in\{{\rm G},{\rm Y},{\rm SK}\}$), we use an explicit formula of the quantity $\frac{|R(\varphi^{0}_{\pi},E,\Lambda)|^{2}}{(\varphi_{\pi}^{0}|\varphi_{\pi}^{0})_{L^{2}}}$ first proved by Dickson, Pitale, Saha, and Schmidt ([7, Theorems 1.13 and 3.13]) when $\Lambda$ is unramified and $l_{1}=l_{2}$ under the refined GGP conjecture for Bessel periods posed by Liu([24]). Owing to a theorem by Furusawa and Morimoto [11, Theorem 8.1], the formula in [7, Theorems 1.13 and 3.13] is extended to vector-valued forms unconditionally so that it can be applied to (5.11); by Theorem 7.1, we can further extend the formula to any ramified $\Lambda$ with square-free conductor.
@@
Let the notations and the assumptions be as in Theorem 5.2.
@@
    ×$\times{\mathcal{M}}^{-8}M^{s-6}\zeta_{M}(1)\zeta_{{\mathcal{M}}M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})$
@@
    where $S(N)$ is the set of all the prime divisors of $N$, and ${\mathcal{A}I}(\Lambda^{-1})$ is the automorphic induction to ${\bf GL}_{2}({\mathbb{A}})$ from the character $\Lambda^{-1}$ of ${\mathbb{A}}_{E}^{\times}/E^{\times}$.
@@
    ×$\times{\mathcal{M}}^{-8}M^{s-6}\zeta_{M}(1)\zeta_{{\mathcal{M}}M}(4)\tilde{\mu}(2D)G(\tilde{\mu})N^{s-1}\tilde{\mu}^{-1}(N)\prod_{p\in S(N)}(1+p^{-1})$
@@
The equation (i) and (ii) is a direct corollary to Theorem 5.2 and [10, Theorem 8.1](for the scalar case we refer to [7, Theorems 1.13 and 3.14]), extended to the ramified $\Lambda$ by Theorem 7.1; note that the polynomial $Q_{S,\varrho}$ (with $S=T_{\theta}$) in [10, (8.2.16)] equals $(-1)^{\frac{l_{1}-l_{2}}{2}}(\frac{2}{\sqrt{|D|}})^{\frac{l_{1}+l_{2}}{2}}\mathsf{v}_{\varrho}^{\theta}$, our $R(\varphi,E,\Lambda)$ equals the quantity $w_{D}(\frac{\sqrt{|D|}}{2})^{-\frac{l_{1}+l_{2}}{2}}\,\mathcal{B}_{\Lambda}(\varphi;E)/(Q_{S,\varrho},Q_{S,\varrho})_{l_{1}-l_{2}}$ defined by [10, (8.2.17)], and $(Q_{S,\varrho},Q_{S,\varrho})_{l_{1}-l_{2}}=(\frac{|D|}{4})^{-\frac{l_{1}+l_{2}}{2}}(\mathsf{v}_{\varrho}^{0},\mathsf{v}_{\varrho}^{0})_{l_{1}-l_{2}}=(\frac{|D|}{4})^{-\frac{l_{1}+l_{2}}{2}}(\mathsf{v}_{\varrho}^{0}\mid\mathsf{v}_{\varrho}^{0})_{\varrho}$. We also note the formula
@@
where $\pi=Y(\pi_{1},\pi_{2})$. This is obtained immediately by comparing the local factors. It is noted in [7, p.296] that each local representation $\pi_{p}$ of $\pi\cong\otimes_{p}\pi_{p}\in\Pi_{\rm cusp}^{(T_{\theta},\Lambda)}(\lambda,N)^{\rm Y,new}$ for $p\mid N$ is of type VIb. Finally, we check the statement (iii). As quoted in [7, Theorem 3.11], the $(T_{\theta},\Lambda)$-Bessel periods of the Saito-Kurokawa lifts are zero unless $\Lambda$ is trivial due to Qiu. By comparing the local $L$-functions ([40, Theorem 5.2(ii)], cf. [30, Theorem 3.1]),
@@
By (5.4), we have that ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})-{\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm G,new}$ is the sum of ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm Y,new}$, ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm SK,new}$ and ${\mathbb{I}}^{(s)}(\lambda,N,f_{S})^{\rm old}$. Our main focus is their values at $s=0$. For those, we have the following upper bounds in the level aspect. For $f_{S}\in{\mathscr{H}}_{S}$, let $\|f_{S}\|_{1}:=\int_{\mathsf{G}({\mathbb{Q}}_{S})}|f_{S}(g_{S})|\,{\rm{d}}g_{S}$ denote its $L^{1}$-norm.
@@
    Let $\lambda$ be fixed as before and $N$ an inert prime in $E/{\mathbb{Q}}$. Then, there exist constants $C_{D,\lambda,\mu},C_{D,\lambda,\mu}^{\prime}>0$ depending only on $D$, $l$, $\mu$, and $f_{S}$ such that
    (5.14) $|{\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm old}}|$ $<C_{D,\lambda,\mu}\|f_{S}||_{1}N^{-8},$
    (5.15) $|{\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm Y,new}}|$ $<C_{D,\lambda,\mu}^{\prime}\|f_{S}\|_{1}N^{-\frac{3}{2}}.$
@@
    Let $N$ be $1$ or an inert prime in $E/{\mathbb{Q}}$. When $\mu\neq{\bf 1}$, ${\mathbb{I}}^{(0)}(l,N,f_{S})^{{\rm SK,new}}=0.$ When $\mu={\bf 1}$, there exists a constant $C>0$ such that
    (5.16) $|{\mathbb{I}}^{(0)}((l,l),N,f_{S})^{{\rm SK,new}}|<C|D|^{\frac{l}{2}}l^{\frac{1}{2}}\|f_{S}\|_{1}N^{-\frac{3}{2}}.$
@@
We have the inequality $|\widehat{f}_{S}(\pi_{S})|\leqslant\|f_{S}\|_{1}$ for all irreducible unitary representation $\pi_{S}$ of $\mathsf{G}({\mathbb{Q}}_{S})$. Because $N$ is a prime, one can check that the summation range of ${\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{{\rm old}}$ is $\Pi_{\rm cusp}^{(T_{\theta},\chi)}(l,N)^{{\rm old}}=\Pi_{\rm cusp}^{(T_{\theta},\chi)}(\lambda,1)$, which is independent of $N$. From the formula of ${\mathfrak{t}}(\pi_{p},\mu_{p})$ in (5.12) for $p=N$, combined with the temperedness of $\pi_{N}$ due to [49] and the matrix of $T_{1,0}$ in [34, Table 3], we get ${\mathfrak{t}}(\pi_{N},\mu_{N})=O(N^{-4})$. This and the equation $[{\bf K}_{\bf f}:{\bf K}_{0}(N)]=N^{3}(1+N^{-2})(1+N^{-1})$ for prime $N$ yield the bound (5.14).